Thursday 29 October 2015

Maintenance Update (painting, lapping, braking)

Wednesday 21st
John T reports:
" I'll admit it, since next time you are in the TPO you will suspect it, that I was at Toddington today. I managed to clean 6 chairs, paint 8 bottoms and cut some bolts.
I felt I owed a visit, especially as I can't attend on Saturday next."

Saturday 24th
Fred, Dixie and Gil passed through Todders, but went on to Winchcombe to play in the siphon.  Bruce painted the rail chairs that John had prepared.  We're still a bit stymied regarding doing any work on the loco while she is on stand-by.

Wednesday 28th
There are three new issues raised (probably while on the Friday Fire & Drive):
58 [MY]: Steam heat safety valve blowing @ 30 psi.
74 [MY]: Handbrake 8-9 turns; needs adjusting.
75 [MY]: LH clack passing to pep pipe live when off.  {Sorry if I misread the shorthand!}

There is a valve on the steam heating pipe at the back of the tender, which is designed to prevent excessive pressure passing into the coaching stock.  Generally, GWSR limits this to 40 psi.  Bruce decided to tackle this.  He extracted its innards and lapped-in the internal valve, and then reassembled it.  Of course, there is no way of verifying the pressure-relief point when there's no pressure, so Bruce just tightened it a shade, and we'll see what happens when we next have steam up.

Bruce was going to help Gil with the tender brakes, but I arrived and took over.  There's a "bottle screw" under the tender that has to be turned to shorten the brake linkages.  It requires the use of a BIG spanner.  It take you all your effort to lift the spanner, let alone turn it!  Nevertheless, Gil & I managed to rotate the bottle-screw one whole turn, which reduced the number of turns on the handbrake to five (to apply the brakes).

Dixie passed by, having already been at Winchcombe.  He picked up a couple of electrical sockets for the siphon; sanded ten brushes for boot scrapers; and then buzzed off again.  Bill, Geof and Ray were all working on the siphon at Winchcombe.

Gil & I sorted through a pile of technical drawings, trying to put them into some sort of order and file them in Carpo's drawers.  A few of these drawings are not on the GWSR system, so Gil has taken them to get them scanned.

Finally, I applied stain to the brushes that Dixie had sanded.

2807 is on stand-by over this weekend.  Next Wednesday and also Saturday, we have to unbox the boiler and give it a wash-out.  The Boiler Inspector comes on Monday 9th.  Wednesday will be boxing-up the boiler again, and lighting a warming fire.  Bruce wants to get the safety valve off and lap that in before the steam test … on Thursday (12th).  We may then be in use on Race Trains that Friday & Saturday.  THEN we should be free to begin winter maintenance!


Sunday 18 October 2015

Maintenance Update (reverser, shackle, meeting)

Saturday 10th
Bruce checked the issues log, and there were no issues from Friday's Fire & Drive.  So, there was not a lot to do.  However, Bruce did remind me that various crews had felt some sort of knocking from the loco.  I had heard a knock once, when walking along side as she was being shunted in the yard.  So, John T and I decided to clamber over her looking for "shiny bits" (i.e. where something might have been knocking or rubbing against something that it shouldn't).

Despite a good crawl inside & out, the shiniest thing we found was a rubbing wear on the right-hand lifting link.  It is clear that the bolt through the forward eccentric rod does catch on the lifting link.  However, there appeared to be a gap of about 1/8 inch.  We waggled the reverser; we heaved on the rods; but we couldn't move anything significantly.  There is a gap at each end of the reversing shaft, where there is a collar to limit the sideways movement.  Could this possibly move enough to cause the bolt & link to collide?  Bruce didn't think so.  Maybe it's just one of those things - when hurtling along at 25 mph, things oscillate and sometimes do 'high fives' with one another?  Either ways, when she's near to a power supply, we'll grind a bit off the end of the bolt.

We think that there is a need to walk along side (as I had done previously) and listen to where this knock is coming from.

Eleanor (not my daughter) had cleaned out the smokebox and the grate.  It was too hot to get inside the firebox and clean the grate properly.  The Wednesday gang will have to do that.

Bruce was giving Eleanor a lesson in vacuums as I clambered into the cab.  I distracted him, and we took a look at the copper pipe with a crack in it.  There was no pressure on the 'clock', though some steam would pull through the whistle.  Guess what?  There were drops of water oozing out of the crack in that pipe!  This despite that coil being closed off.  The consensus view is that the packing in the Y-splitter expands/contracts dependent upon steam pressure, and under little pressure it allows steam to squeeze past the on/off cock.  Once pressure rises, it seals.

Bruce investigated making a new pipe.  There is no point in trying to silver-solder the outside of a pipe that is under boiler pressure.  The ends of the pipe are formed (swaged).  We have the technology, but not the pipe.  I asked Gil to order some 1/2 inch thick-walled copper pipe.  He things it comes in ten-foot lengths.

In odd moments, John cleaned up a couple of rail chairs and chopped the bolts of many more.  I applied a top coat to 7 in the production line, and then did some tidying up of the pile of rail chairs awaiting attention.

Fred, Gil, Geof and Bill were all working in the siphon.  Geof & Gil were forming the joints (tenons) in the new door frame.  Bill was painting the metalwork of the ceiling.  Fred was doing something on the outside!  I was only calling in to get Gil's autograph on the cheque for our 100 Club winners.  But guess what I spotted … two LMS chairs and one LNER ….

Sunday 11th
… which are now in the pile at Todders.

Down there on crossing duty at the Diseasel Gala, I realised that (yet again) no one had put the cap on the chimney!  So, I did, of course.

Wednesday 14th
There was only Bruce & myself today.  We decided to try to fix the bolt (strictly, it's a tapered pin with a castellated nut on the end) that is making the shiny mark on the hanging link.  Bruce connected up the power - we had to run the 110V from the end of the pit down to the loco, which was well past the pit.  Then he read the notice that says not to use the 110V in road 8, so he had to move it to road 9's pit.  Finally, after covering everything with cloths, I was able to angle-grind a smidgeon off the end.  Bruce, being today's gopher, brought the red paint, and I applied a little to the end of the bolt and to the hanging link.  We'll find out on Saturday (after Friday's Fire & Drive session) if they still have a spatial coincidence problem.

Gil arrived, partly to deliver an invoice to me and partly to bring the newly-purchased adjustable reamer … and then discuss with Bruce the plan of attack.

Meanwhile, I painted lettering on seven boot scrapers and fitted a brush to each of six.

Gil buzzed off to Winchcombe, so Bruce & I decided to wend our ways, too.  Bruce went home … to do some homework; I boxed up the six boot scrapers and adjourned.

Today we had to say 'farewell' to a faithful friend, who has faultlessly carried out its duties for some 30-odd years.  Someone noticed that the shackle that has helped heave many an item up, in & out of 2807 is marked "for recovery use only", so we are banned from using it to lift anything!  So? … We were only "recovering" the valve rods from the cylinders …

Saturday 17th
Today was a 2807 Board Meeting (which means that little work gets done!).

Bruce & David struggled past us in order to get some jobs done.  They began by preparing and welding the seam of the toolboxes on the tender.  These had been coming apart for a while.  Having been spotted wielding a welder, David was then co-opted into doing some welding for Dinmore Manor!

Bruce move on to removing the flange from the out-flow side of our blow-down valve.  JC had asked to borrow this in order to make a pattern for it.

There will be a period of silence on the 2807 updates, because I am otherwise engaged for a bit.  Rest assured that I have not forgotten about you!


Friday 9 October 2015

Maintenance Update (condenser, copper, breather)

Thursday 1st Oct
I popped in to Todders just to finish off the boot scraper ordered by Dame Janet Trotter.  John P [Loco Dept] was lighting up 2807 ready for the Fire & Drive on Friday.  We had a chat about the oil level in the RH big end, and he said he'd keep watch.

2807 Fireman's report from Friday 2nd Oct.
"Right Hand Connecting rod big end oil reservoir.
I spoke to Steve Oddy about the supposed problem prior to leaving the shed. We checked the reservoir at Cheltenham on the first trip and again at Toddington when stopped for lunch. There had been no excessive oil loss and the connecting rod temperature was normal i.e. slightly warmer than the coupling rod journals. This was the situation throughout the remaining Fire and Drive service. I checked the Connecting Rod oil reservoir during disposal and found that the level had changed very little. The big end temperature had also remained as normal. Looking through the drivers report cards I suspect a faulty cork had caused the oil loss problem.

Steam Condenser Coil.
This was interesting as it behaved exactly as it had done the previous time I reported it when on duty with Steve. In that case, as with on Friday the engine was "cold" and had not been in service the previous day. Anyway, overnight the boiler pressure was about 10psi from my warming fire. As the pressure reached 40-50psi a well-defined steam leak on the first bend in that coil could be seen. As boiler pressure increased further to normal working pressure the leak slowly disappeared and remained so throughout the day. I suspect the "T" cock is passing a little when cold. The lubricator working on the other coil functioned normally throughout the day."

Saturday 3rd
Bruce checked the oil level and it was roughly at the top of the restriction pipe (i.e. still plenty in the reservoir below it).

No new issues logged from the Friday outing.

Gil was here briefly, and painted a few bottoms…  Thereafter he scurried off to Winchcombe (as is his wont).

Bruce had been concerned about the lack of lock-nuts on the loco's vacuum reservoir (beneath the cab).  The studs are a tad short, but Bruce made one lock nut and fitted it.  There has been no hint of the nuts working loose.

Later, Bruce was diverted onto a 5542 problem.  Someone (not knowing their own strength) had broken part of its steam heating valve.  It had been silver-soldered back together, and Bruce got the job of cleaning up the carrot (i.e. the tapered internal plug).  Although he made a good job of this, when the valve was assembled and tested, the repair fell apart (at which point Bruce handed it back!).

As we have 2 orders plus a potential order for ten more boot scrapers, Dixie and I set to on cleaning some.  By 3.30 pm we were exhausted, having prepared 13 rail chairs.  Gil painted the first four's bottoms before dashing off to Winchcombe to play with the siphon.  Dixie finished them off, so we have 13 shiny black bottoms pointing skywards.

Sunday 4th
Bruce (who lives a stone's throw from the station) heard a familiar whistle!  It transpires that the 5542 had a bit of an issue: "Cab floor broken in two places", hence it was "red-carded" and 2807 came to the rescue.  This explained the puzzle of Malcolm R's photo of 2807 on the Sunday service!

Monday 5th
I popped in to complete a couple of boot scraper orders.  I checked the oil in the big end, and it was fine - not even dropped down to the level of the top of the restrictor pipe.  I also fitted the chimney cap - we'll have to have a word with the Loco Dept.  That's the 2nd time I've had to fit the chimney cap (that the crew should do at the end of the day).  Meanwhile, it had rained during the night (and the cap is partly there to prevent rain ingress to the smokebox, where ash + rain = acid).

Wednesday 7th
Gil & Bruce discussed plans for the winter maintenance … then Gil disappeared to Winchcombe (again).  Bruce then tightened up the balance weights on the driving wheels.  They are lead, and do work loose over time.  It's easy enough to threaten them with a hammer and force them to tighten.  They are not loose enough to fall off!

John G painted about 10 chair tops in Deproma (primer/undercoat); David M joined us after lunch and completed the final few.  I, too, worked on boot scrapers for the day (surprise, surprise).

During the afternoon, Bruce removed the section of copper pipe between Y-splitter and coil to see if we could see a split in it.  People have been recording issues about it since July.  Yes indeed, there are signs of splits on the rather sharp bend where it turns to fit onto the Y-splitter.

Following on from the above, there was one new (!) issue recorded in the crew's log:

73 [AM]: Suspect split in LH condensing pipe, where it enters T cock.  {This is the same as issues 54 and 47, and that Bruce verified today.  We can do nothing now; it will need a new pipe making.}

Thursday 8th
I spoke with SO today about our big end problem.  He's on 2807 for the Fire & Drive on Friday.  It seems there was a misunderstanding (a mild ambiguity, in fact).  It was he who discovered the cork with no breather in it, which he replaced … and the oil-loss problem went away.  The full fact is that the cork had a hole in it where there should have been a cane breather!  No one had noticed!  That's why it was losing oil.  We assumed that the issue 72 report "cork had no core" meant that it was a solid cork … but no!  The core had fallen out.  Doh !!!


Thursday 1 October 2015

Maintenance Update (felt, small end, nut)

Saturday 26th
This was our AGM day.  It was good to see so many (mostly familiar) faces.  After the AGM, many of us took a trip behind 2807.

There was some discussion (in the TPO) about the loss of oil from the RH big end.  It really is a puzzle.  The oil pot has a vertical pipe within it that has a restriction through it.  Oil should splash around as the loco moves, and some will go down the pipe.  At the bottom of the pipe is a felt pad.  This should restrict the flow of oil and make the pot full last all day.  So, why doesn't it?  The decision was to remove the con rod on Wednesday and have a look at the pad.

I did suggest that replacing the cork (a plug in the top of the oil pot, which has a cane breather through the middle) with a solid one should create a partial vacuum and thereby restrict the oil flow.  This purely as a temporary measure until we could remove the rod.

Monday 28th
I went over to Todders to do some chair painting - we have two or three boot scraper orders that need progressing.  2807 was in the yard, and the RH side con rod was positioned exactly at the angle that it has to be when being removed.  The position is fairly critical, because the step gets in the way, and also the small end has to be backwards of the slide bar bracket.  This was either a stroke of luck, or a deliberate act by the driver - who knows?!  So, I took the decision to fit a Not To Be Moved board on the loco!

In this low position, the oil pot is easily accessible, so I filled it up.  Let's see how much runs out over the two day period.  The oil level can be expected to fall to that of the top of the pipe (but no motion = no splashing, so the level should not drop below that … unless there's a serious problem in the pot!).

I checked the Issues Log for new reported issues from Saturday.  Two:

57 [JC] LH lifting link top pin loose.  Pin tightened & new split pin fitted.  {Thank you, JC}

72 [anon] @ cork had no core.  Refitted cork.  Oil usage decreased.  Continue to monitor.  {so much for the solid cork test - driver spotted it and replaced it with a breathable one!}

The numbering is up the creek .. it was young Jon W's fault!  ;-)

Wednesday 30th
Gil, Bruce, John G & I (aided from time to time by Loco Dept chaps) tackled the removal of the RH con rod.  The oil level had dropped to the top of the restrictor tube (as one would have expected), so that gave no clue about why it loses so much so quickly.  It transpired that the wheels were not 100% perfectly lined up … we had to push the loco back by about 2".  This wasn't too difficult, once the three of us were pushing in the same direction.

The vacuum pump has to be disconnected first (because it is linked to the cross-head).  The big end came off easily.  The little end is always a challenge, and needs an extractor.  So, it took Bruce & Gil a bit longer to get that free.

Sliding the rod off was easy.  We then turned it over to be able to see the felt pad.  It was in place and intact.  We could see nothing wrong at all.  Nevertheless, it was decided to replace the felt in both ends.  I ran a wire through the constrictor, just to be sure it was not blocked.  It was fine.

Today was 3850's last day in service before its Heavy General Overhaul, and John G took the opportunity to ride behind as far as Winchcombe.  He was able to report that Fred, Ray and Bill were working on the siphon restoration.  They had completed the construction of a new pair of doors and were just making minor adjustments to its fit.  They have the wood ready for constructing the other doors, but will test this one in each place first, just to verify that the doors are all exactly the same size!

Bruce cut new felt pads.  These were inserted and then the con rod was fitted back.  Once again, it was the small end that caused trouble.  Getting the castellated nut lined up with the split pin hole took for ever.  It was either the nut being too slack, or too tight.  Some fiddling of the washer, nut, its castellations and the split pin finally got it all to fit together.

We are none the wiser as to why the oil should have run out so quickly (i.e. after each single trip), unless … did the driver realise that the level will fall rapidly to the level of the top of the constrictor, and then very slowly thereafter?  Maybe he thought that the pot always needed to be full above the constrictor top?  Who knows?!

During a couple of slack periods, John G and I managed to restock the Flag & Whistle (they having sold five boot scrapers) and do a little more on the production line.

2807 is on Fire & Drive duty on Fridays during October.  She's not rostered to be in public service now until next year.